### v1

# XSNAP: An X-ray Supernova Analysis Pipeline with Application to the Type II Supernova 2024ggi

Abstract
X-ray observations of Type II supernovae (SNe II) probe the physics of supernova (SN) shocks and the mass-loss histories of their progenitor stars. We present multi-epoch, X-ray observations of SN II 2024ggi ($D\approx 7.2$ Mpc) from Swift-XRT, Chandra X-ray Observatory and XMM-Newton, which cover $\sim 1-344$ days since first light. We analyze these observations using a new open-source Python package called XSNAP, which standardizes a unified command-line interface for instrument-specific reduction and spectral extraction. XSNAP introduces application programming interfaces for per-epoch spectral modeling through PyXspec and emcee Markov chain Monte Carlo fitting. We employ XSNAP to model the multi-epoch X-ray spectra of SN 2024ggi with an absorbed thermal bremsstrahlung model and calculate a steady progenitor mass-loss rate of $(6.2\pm 0.2)\times 10^{-5}\,M_{\odot}\,\mathrm{yr^{-1}}$ $(v_{\rm wind}=20\ \rm km\ s^{-1})$, for which the detected X-ray emission traces the final 117 years before explosion. The software is publicly available on GitHub, with a released package on the Python Package Index (PyPI).
[…]
facilities: CXO, Swift (XRT), XMM-Newtonsoftware: Astropy (Astropy Collaboration et al., 2013, 2018, 2022), CIAO (A. Fruscione et al., 2006), corner (D. Foreman-Mackey, 2016), emcee (D. Foreman-Mackey et al., 2013), HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014)Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), Heasarc), NumPy (C. R. Harris et al., 2020), Pandas (Wes McKinney, 2010; T. pandas development team, 2020), PyXspec (C. Gordon & K. Arnaud, 2021), SAOImage DS9 (W. A. Joye & E. Mandel, 2003), SAS (SAS development team, 2014), SciPy (P. Virtanen et al., 2020), XSPEC (K. A. Arnaud, 1996)
show]ff10@illinois.edu
1 Introduction
X-ray observations of supernovae (SNe) have become increasingly common in the last few decades. Notably, improvements in X-ray spectroscopy allow for the study of element distributions and temperatures in supernova remnants as well as distinguish between thermal (e.g., Bremsstrahlung) and non-thermal (e.g., synchrotron or inverse Compton) emission (J. Vink, 2012). A recent census from V. V. Dwarkadas (2025) reports that almost all SNe detected in X-rays are from a core-collapse origin except for SN 2012ca which was classified as Type Ia-CSM by O. D. Fox et al. (2014). V. V. Dwarkadas (2025) suggests the following X-ray characteristics of core-collapse SN subtypes: Type Ib/c SNe are typically non-thermal; Type IIP with low progenitor mass-loss $(\dot{M}\sim 10^{-7}\rm\ M_{\odot}\ yr^{-1})$ are likely non-thermal, with thermal emission becoming dominant as $\dot{M}$ increases; Type IIn events are the most X-ray luminous and clearly thermal; and Type IIb SNe are mostly thermal with a subclass of compact progenitors (cIIb) showing non-thermal emission.
X-ray emission in SNe arises when shocks, created from the fast SN ejecta interacting with the circumstellar medium (CSM), sweep up the CSM, heat the gas to high temperatures, and then the post-shock region cools. In mass coordinates, the resulting shock waves propagate in two directions: inward (reverse shock) and outward (forward shock). Radiation from the reverse shock typically has lower temperatures than that from the forward shock (R. A. Chevalier & C. Fransson, 2017). Moreover, X-rays provide a direct probe of the CSM density profile, enabling estimates of the progenitor’s mass-loss rate in the final years before explosion (R. A. Chevalier, 1982a, b; R. A. Chevalier & C. Fransson, 2017; V. V. Dwarkadas, 2025). For example, density profiles inferred from X-ray modeling of Type II SNe are often compared to the self-similar solution of R. A. Chevalier (1982b), in which $\rho\propto r^{-s}$ with $s=2$ for a typical Type II SNe.
Observations over the past two decades have validated and revealed diverse results across SN subtypes. Type Ib/c SNe usually show low mass-loss rates, except some with thermal emission whose mass-loss may reach $\gtrsim 10^{-5}\rm\ M_{\odot}\ yr^{-1}$, with most extreme cases can be up to $10^{-3}\rm\ M_{\odot}\ yr^{-1}$ (V. V. Dwarkadas, 2025). Type IIP/L SNe typically exhibit lower mass-loss rates of around $\sim 10^{-7}-10^{-5}\rm\ M_{\odot}\ yr^{-1}$, e.g., SN 2013ej (S. Chakraborti et al., 2016), SN 2017eaw (T. Szalai et al., 2019), and SN 2004et (K. Misra et al., 2007) (although there are exceptions e.g., SN 2023ixf, A. J. Nayana et al., 2025). Type IIn SNe have the greatest mass-loss rates among SN subtypes, i.e. $\gtrsim 10^{-3}\rm\ M_{\odot}\ yr^{-1}$, e.g. SN 2017hcc (P. Chandra et al., 2022), SN 2010jl (C. Fransson et al., 2014), SN 2020wyx (R. Baer-Way et al., 2025), or SN 2005ip (S. Katsuda et al., 2014). Finally, Type IIb SNe generally have mass-loss rates of around $10^{-5}-10^{-3}\rm\ M_{\odot}\ yr^{-1}$, e.g., SN 2018gk (S. Bose et al., 2021), SN 2013df (K. Maeda et al., 2015), and SN 2008ax (P. W. A. Roming et al., 2009).
An example of a well-studied Type IIP SN in the X-rays is SN 2023ixf, discovered by K. Itagaki on 2023 May 19 in the M101 galaxy (K. Itagaki, 2023). It was first observed in at X-ray wavelengths with the Swift X-ray Telescope (XRT) at $\sim 1$ day after first light, which was then complimented with follow-up observations from the Chandra X-ray Observatory (CXO), XMM-Newton, and Nuclear Spectroscopic Telescope Array (NuSTAR) (W. Jacobson-Galán, 2025). Multiple analyses have been done with the absorbed plasma model (P. Chandra et al., 2024a) and the absorbed Bremsstrahlung model (B. W. Grefenstette et al., 2023; E. A. Zimmerman et al., 2024; S. Panjkov et al., 2024; A. J. Nayana et al., 2025; W. V. Jacobson-Galán et al., 2025a) across the soft and hard X-rays, with the X-ray luminosity of $\sim 10^{39}-10^{40}\ \rm erg\ s^{-1}$ within the first $\sim 10^{2}\ \rm days$. The findings of these studies are consistent and imply that SN 2023ixf interacted with dense, confined CSM and the red supergiant (RSG) progenitor mass-loss rate was around $10^{-4}-10^{-3}\ \rm M_{\odot}\ yr^{-1}$, one to two order magnitude higher than the upper-limit of SNe IIP detected in X-rays (V. V. Dwarkadas, 2025).
SN 2024ggi, another nearby Type II SN (W. Hoogendam et al., 2024; Q. Zhai et al., 2024), is a well-monitored SN across multiple wavelengths: ($\gamma$-rays (G. Marti-Devesa & Fermi-LAT Collaboration, 2024), X-rays (J. Zhang et al., 2024a; R. Margutti & B. Grefenstette, 2024), optical bands (T. W. Chen et al., 2024; T. Killestein et al., 2024; B. Kumar et al., 2024; F. D. Romanov, 2024; L. Wyrzykowski et al., 2025), radio frequencies (S. Ryder et al., 2024), and centimeter wavelengths (P. Chandra et al., 2024b)). The SN was first discovered by the Asteroid Terrestrial-impact Last Alert System (ATLAS) on 2024-04-11 (MJD 60411.14) (J. Tonry et al., 2024; S. Srivastav et al., 2024; T. W. Chen et al., 2025) and was officially classified as a Type IIP (J. Zhang et al., 2024b; T. W. Chen et al., 2025). Previous studies have analyzed UV/optical/IR observations and inferred that the RSG progenitor mass-loss rate was $10^{-3}-10^{-2}\ \rm M_{\odot}\ yr^{-1}$ (M. Shrestha et al., 2024; J. Zhang et al., 2024b; W. V. Jacobson-Galán et al., 2024a; A. Aryan et al., 2025; T. W. Chen et al., 2025; K. Ertini et al., 2025). Additionally, D. Xiang et al. (2024) estimated the progenitor mass-loss rate is $3\times 10^{-6}\ \rm M_{\odot}\ yr^{-1}$ from pre-explosion imaging and suggested that SN 2024ggi had an enhancement of mass-loss rate in the century leading up to the explosion. Use of nebular spectroscopy has indicated a progenitor mass in the range of $11-15.2\ \rm M_{\odot}$ (D. Xiang et al., 2024; A. Aryan et al., 2025; L. Dessart et al., 2025; K. Ertini et al., 2025; L. Ferrari et al., 2025; E. Hueichapán et al., 2025). In this work, we present and analyze the X-ray observations of SN 2024ggi from Chandra X-ray Observatory, Swift X-ray Telescope, and XMM-Newton. SN 2024ggi is located at $(\alpha=\rm 11^{h}18^{m}22.09^{s},\delta=-32^{\circ}50^{\prime}15.26^{\prime\prime})$ in the host galaxy NGC 3621. We adopt a time of first light corresponding to MJD $60410.80\pm 0.34$ days, redshift of $z=0.002215$ (W. V. Jacobson-Galán et al., 2024a), and a distance of $7.2\pm 0.2$ Mpc (A. Saha et al., 2006).
We present multi-epoch X-ray observations of SN 2024ggi in Section 2. Our analysis of SN 2024ggi and its surrounding CSM are described in Section 3. In Section 4, we introduce the pipeline we have developed, XSNAP, and its design and availability. We discuss our findings in Section 5 and summarize them in Section 6.
2 Observations
Table 1: Best‐fitting Parameters of the Absorbed Thermal Bremsstrahlung Model for SN2024ggi
[…]
(days) $\left(10^{22}\ \rm cm^{-2}\right)$ (keV) (Absorbed) (Unabsorbed) ($10^{-4}$)
3.33 Swift-XRT $7.36^{+3.93}_{-2.89}$ 47.81 $12.57^{+0.13}_{-0.14}$ $12.30^{+0.13}_{-0.14}$ $0.96^{+0.41}_{-0.30}$
8.36 Swift-XRT $9.35^{+6.33}_{-3.92}$ 37.97 $12.54^{+0.12}_{-0.12}$ $12.23^{+0.12}_{-0.12}$ $1.09^{+0.56}_{-0.37}$
10.90 CXO $2.83^{+0.57}_{-0.51}$ 35.54 $12.51^{+0.04}_{-0.04}$ $12.32^{+0.04}_{-0.04}$ $0.88^{+0.10}_{-0.09}$
16.20 CXO $1.33^{+0.42}_{-0.38}$ 32.18 $12.69^{+0.05}_{-0.04}$ $12.56^{+0.05}_{-0.04}$ $0.51^{+0.07}_{-0.06}$
24.49 Swift‐XRT $0.06^{+0.53}_{-0.06}$ 29.02 $12.81^{+0.12}_{-0.17}$ $12.77^{+0.12}_{-0.17}$ $0.31^{+0.13}_{-0.08}$
34.76 Swift‐XRT $0.00^{+0.37}_{-0.00}$ 26.59 $12.87^{+0.14}_{-0.16}$ $12.84^{+0.14}_{-0.16}$ $0.26^{+0.11}_{-0.08}$
43.20 Swift‐XRT $0.00^{+0.08}_{-0.00}$ 25.18 $12.96^{+0.13}_{-0.18}$ $12.94^{+0.13}_{-0.18}$ $0.21^{+0.07}_{-0.05}$
55.03 XMM $0.00^{+0.02}_{-0.00}$ 23.7 $13.10^{+0.06}_{-0.08}$ $13.07^{+0.06}_{-0.08}$ $0.15^{+0.02}_{-0.02}$
85.40 XMM $0.01^{+0.05}_{-0.01}$ 21.24 $13.25^{+0.04}_{-0.05}$ $13.22^{+0.04}_{-0.05}$ $0.11^{+0.01}_{-0.01}$
Note. — All errors are reported at $1\sigma$ c.l.
Figure 1: Top: Plot of fitted spectrum from CXO at $\delta t=10.90\ \rm days$ (left) and $\delta t=16.20\ \rm days$ (right) since first light. Blue (left) and green (right) points are observed X-ray spectra and black dashed lines are the best-fit absorbed thermal Bremsstrahlung model. Bottom: Plot of fitted spectrum from XMM data $\delta t=55.03\ \rm days$ (left) and $\delta t=85.40\ \rm days$ (right) since first light. Observed spectra are represented in blue (EPIC-pn), orange (EPIC-MOS1), and magenta (EPIC-MOS2) and black dashed lines are the best-fit absorbed thermal Bremsstrahlung model.
Figure 2: Left: Plot of fitted unabsorbed $0.3-10\ \rm keV$ X-ray luminosity light curve from XMM-Newton (red circles), CXO (cyan squares), and Swift-XRT (orange triangles) observations. The red dashed line is the fitted line with $L\propto t^{-0.99}$. Right: Plot of unabsorbed $0.3-10\ \rm keV$ X-ray luminosity light curve of SN 2024ggi compared to a sample of Type IIP SNe. References: E. M. Schlegel (1999, 2001), D. Pooley et al. (2002), D. Pooley & W. H. G. Lewin (2002, 2004), K. Misra et al. (2007), S. Immler et al. (2007), S. Immler & P. J. Brown (2012), S. Chakraborti et al. (2013, 2016), T. Szalai et al. (2019), I. Irani et al. (2024), A. J. Nayana et al. (2025)
Figure 3: Left: Plot of fitted density profile. The blue circles are the data derived from analysis and the red dashed line is the fitted line with $\rho\propto r^{-2}$. Right: Plot of fitted density profile to a previous study by W. V. Jacobson-Galán et al. (2024a). The r1w4 (green dashed line), r1w6 (red dashed line), and m1em2 (orange dashed line) are best-matched models from CMFGEN model grid (W. V. Jacobson-Galán et al., 2024a, b). The magenta star is the transition point where the CSM goes from optically thick to thin to electron scattering (W. V. Jacobson-Galán et al., 2024a). The blue circles are the data derived from analysis and the blue dashed line is the fitted line with $\rho\propto r^{-2}$. The purple pentagons are SN 2023ixf densities derived from X-ray observations (A. J. Nayana et al., 2025). The grey dashed lines are reference density profile lines if the mass-loss rates are $10^{-4}\ \rm M_{\odot}\ yr^{-1}$, $10^{-5}\ \rm M_{\odot}\ yr^{-1}$, and $10^{-6}\ \rm M_{\odot}\ yr^{-1}$ respectively for $v_{\rm wind}=20\ \rm km\ s^{-1}$.
2.1 Chandra X-ray Observatory
The Chandra X-ray Observatory (CXO) ACIS-S observed SN 2024ggi in two epochs: one on 2024 April 21 (ObsID: 29383; $\delta t=10.90\ \rm days$; PI: E. Zimmerman) with an exposure time of 14.88 ks and the other on 2024 April 26 (ObsID: 29384; $\delta t=16.20\ \rm days$; PI: E. Zimmerman) with an exposure time of 14.58 ks. Details of these two observations are reported in Table 3.
We reduced both observations using CIAO v4.17 (A. Fruscione et al., 2006) and their corresponding calibration files. Source detection and estimating source-free background regions were utilized through ximage v4.5.1 from HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014)Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), Heasarc) with signal-to-noise ratio threshold of three $(\rm SNR_{thresh}=3)$ and human vetting with SAOImage DS9 (W. A. Joye & E. Mandel, 2003). For each epoch, we extracted the spectrum via specextract and used a circular source region with a radius of $2^{\prime\prime}$ and circular background region with a radius of $45^{\prime\prime}$.
2.2 XMM-Newton
Similar to CXO, XMM-Newton (XMM) observed SN 2024ggi in two epochs: one on 2024 June 04 (ObsID: 0882480901; $\delta t=55.03\ \rm days$; PI: W. V. Jacobson-Gálan) and the other on 2024 July 5 (ObsID: 0882481001; $\delta t=85.40\ \rm days$; PI: W. V. Jacobson-Gálan). Details of these two observations are reported in Table 3.
We reduced both observations from the three instruments of XMM, i.e. European Photon Imaging Cameras (EPIC)-pn, MOS1, and MOS2, using the Scientific Analysis System (SAS) v22.1.0 (SAS development team, 2014) and corresponding calibration files. Source detection and estimating source-free background regions were utilized through ximage v4.5.1 from HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014)Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), Heasarc) with signal-to-noise ratio threshold of three $(\rm SNR_{thresh}=3)$ and human vetting with SAOImage DS9 (W. A. Joye & E. Mandel, 2003). For each epoch, we extracted the spectrum via evselect and used a circular source region with a radius of $25^{\prime\prime}$ and circular background region with a radius of $125^{\prime\prime}$.
2.3 Swift X-ray Telescope
The X-Ray Telescope (XRT; D. N. Burrows et al. 2005) onboard the Neil Gehrels Swift Observatory (N. Gehrels et al., 2004) started observing SN 2024ggi on 2024 April 11 UT 14:04:10 $(\delta t=0.79\ \rm days)$. In this work, we analyzed XRT datasets collected until 2025 March 20. As each observation has relatively small exposure time $(\sim 1.5\ \rm ks)$, we binned a few epochs together, ranging from 10-day to 30-day intervals until the source was detected. However, SN 2024ggi was not detected in XRT images after 2024 June 02 $(\delta t=52.20\ \rm days)$ and these observations were analyzed as upper-limits (see Section 3). Details of each stacked observation is reported in Table 3.
We reduced each stacked observation using HEASoft v6.35 (Nasa High Energy Astrophysics Science Archive Research Center (2014)Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), Heasarc), the Swift XRT Data Analysis Software (SWXRTDAS; version 3.7.0), and the Swift Calibration Database (CALDB) version 20250609. Source detection and estimating source-free background regions were utilized through ximage v4.5.1 from HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014)Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), Heasarc) with signal-to-noise ratio threshold of three $(\rm SNR_{thresh}=3)$ and human vetting with SAOImage DS9 (W. A. Joye & E. Mandel, 2003). For each epoch, we extracted the spectrum via xrtpipeline, followed the standard practice from P. A. Evans et al. (2009); R. Margutti et al. (2012), and used a circular source region with a radius of $25^{\prime\prime}$ and circular background region with a radius of $125^{\prime\prime}$.
3 Analysis
3.1 X-ray Spectral Modeling
We modeled and fit each observation in Table 3 with PyXspec (C. Gordon & K. Arnaud, 2021), the Python interface to XSPEC (K. A. Arnaud, 1996). We fit them separately except for XMM-Newton data where we did a joint fit for all three instruments, EPIC-pn, MOS1, and MOS2. In this work, we adopt the solar abundances from M. Asplund et al. (2009), i.e., abund aspl within PyXspec. We set the Galactic neutral hydrogen (NH) column density in the direction of SN2024ggi of $\mathrm{NH}_{\mathrm{MW}}=6.59\cdot 10^{20}\ \mathrm{cm}^{-2}$ (HI4PI Collaboration et al., 2016). For the spectral fits, we restricted the Chandra/ACIS data to the 0.58keV0.5-8\ $\mathrm${keV} energy range, while the Swift-XRT and XMM-Newton spectra were fit in the 0.310keV0.3-10\ $\mathrm${keV} band.
We then fit these data using an absorbed thermal Bremsstrahlung model (tbabs*ztbabs*bremss), but were unable to constrain the model temperature. Therefore, we estimated and fixed the temperatures of each observation using the self-similar solution $T=34\ \rm keV\ (t/13d)^{-0.25}$ (P. Chandra et al., 2024a). This temperature evolution follows the self-similar solutions by R. A. Chevalier (1982b); R. A. Chevalier & C. Fransson (2017) for a standard wind with a constant mass-loss rate, $\rho\propto r^{-2}$. For such a case, the post-shock temperature evolves as $T\propto t^{-2/(n-2)}$, which becomes $T\propto t^{-0.25}$ after assuming $n=10$. We adopted a normalization of $T=34~\rm keV$ at $t=13~\rm d$, as utilized for SN 2023ixf by P. Chandra et al. (2024a). Finally, we successfully modeled the intrinsic NH and Bremsstrahlung normalization and estimated the absorbed and unabsorbed fluxes in $0.3-10\ \rm keV$. We also repeat this analysis while adopting the temperature evolution of SN 2023ixf, $T\propto t^{-0.5}$ (A. J. Nayana et al., 2025), which yields consistent results to those found with a self-similar solution temperature decline rate. Based on the resulting best‐fit parameters, the X-ray spectra of SN 2024ggi reveal SN ejecta interaction with dense CSM: (i) the emission was highly absorbed at early times, (ii) the absorption decreased with time, and (iii) of thermal nature, consistent with previous studies (R. Margutti et al., 2017; B. P. Thomas et al., 2022). The best-fit parameters are reported in Table 1 and contour plots of the first four epochs data from Swift-XRT and CXO are shown in 4. In addition, the fitted spectrum for CXO and XMM-Newton observations are shown in Figure 1.
For observations where SN 2024ggi was not detected, we simulated the spectrum using the FakeIt command from PyXspec and used the faked spectrum flux and count rate ratio to estimate the flux upper-limit in $0.3-10\ \rm keV$. We set the intrinsic NH for non-detections to be $\rm NH_{int}=0$ and did not include their best-fit parameters for further analysis. We converted all the unabsorbed fluxes to luminosities and fit a power-law with emcee Markov chain Monte Carlo (MCMC). We found that the intrinsic $0.3-10\ \rm keV$ luminosity has a decay of $L_{X}\propto t^{-0.99}$ at $\delta t\gtrsim 8.36$ days. Additionally, in Figure 2, we compare the X-ray light curve to a sample of SNe II compiled by I. Irani et al. (2024); A. J. Nayana et al. (2025).
3.2 CSM Density Profile Analysis
We estimated the CSM density profile, $\rho_{\rm CSM}(r)$, by using the normalization of thermal bremsstrahlung model as it is directly proportional to the emission measure (EM). EM is described by:
$EM\equiv\int n_{e}n_{I}\ dV$ (1)
where $n_{e}$ and $n_{I}$ are the electron and ion number densities in the emitting volume $V$. For a thermal bremsstrahlung model in Xspec, the spectral normalization is defined by:
$\rm Bremss_{Norm}\equiv\frac{3.02\times 10^{-15}}{4\pi d^{2}}EM$ (2)
where $n_{e}$ and $n_{I}$ are in $\rm cm^{-3}$ and $d$ is the distance to the source in $\rm cm^{-1}$. We follow the procedure of D. Brethauer et al. (2022) to calculate the unshocked CSM density profile, which goes as:
$\rho_{\rm CSM}(r)=\frac{m_{p}}{4}\left(\frac{2\times\rm EM(r)\mu_{e}\mu_{I}}{V_{\rm FS}(r)}\right)^{1/2}$ (3)
where $\mu_{e}$ and $\mu_{I}$ are the mean molecular weight of electron and ion respectively; $m_{p}$ is the proton mass; $\mathrm{V_{FS}}=\frac{4\pi}{3}f\left(R_{\rm out}^{3}-R_{\rm in}^{3}\right)$ is the emitting forward shock (FS) volume, with $R_{\rm in}$ and $R_{\rm out}$ are the inner and outer radius of the shell respectively and $f$ is the filling factor. In this work, we adopt $R_{\rm out}=1.2R_{\rm in}$, $f=1$, $\mu_{e}=1.14$, and $\mu_{I}=1.24$ (assuming solar values). We also estimated the shock radius by $R_{\rm shock}=v_{\rm shock}\times t$, where $t$ is time since explosion and assumed an average FS velocity $v_{\rm shock}=10^{4}\ \rm km\ s^{-1}$. The mass-loss rate was then estimated by fitting the density profile to $\rho_{CSM}(r)=\frac{\dot{M}}{4\pi r^{2}v_{\rm wind}}$ with the emcee package. Incorporating $v_{\rm wind}=20\ \rm km\ s^{-1}$, we inferred a mass-loss rate of $\dot{M}=(6.2\pm 0.2)\times 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ from the detected emission for the last 117 years before explosion, which lies between the upper-limit of the mass-loss rates of Type IIP SNe $\sim 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ and the lower-limit of the mass-loss rates of Type IIn SNe $\sim 10^{-3}\ \rm M_{\odot}\ yr^{-1}$ (V. V. Dwarkadas, 2025). The density profile and its comparison to SN 2023ixf and density profile models by W. V. Jacobson-Galán et al. (2024a) are shown in Figure 3.
4 The Pipeline
Each X-ray observatory, such as CXO, XMM-Newton, and Swift-XRT, and Nuclear Spectroscopic Telescope Array (NuSTAR), employs mission-specific data formats, calibration pipelines, and analysis software. Consequently, this hinders reproducibility in analyzing X-ray observations as astronomers need to work out distinct guidelines for each mission. While there have been past efforts to streamline the process, such as YAXX [^1] https://cxc.harvard.edu/contrib/yaxx (Yet Another X-ray Xtractor; T. Aldcroft 2006), which supported both CXO and XMM-Newton data, these tools are no longer actively maintained. More recent softwares, like XGA [^2] https://xga.readthedocs.io/en/latest/ (X-ray: Generate and Analyse; D. J. Turner et al. 2022) or BXA [^3] https://johannesbuchner.github.io/BXA/ (Bayesian X-ray Analysis; J. Buchner et al. 2014), offer modern functionality but are limited to a single instrument (e.g., XMM-Newton), not designed as end-to-end pipelines from spectral extraction, or not specifically developed for SNe, even though some may still be adaptable for such studies. At present, there is no open-source, Python-based pipeline that unifies spectral extraction and analysis of SN X-ray spectra from CXO, XMM-Newton, Swift-XRT, and NuSTAR. Consequently, we developed XSNAP (X-ray Supernova Analysis Pipeline), an open-source Python package intended to unify end-to-end X-ray data processes from spectral extraction to CSM analysis.
XSNAP is organized into two stages: spectrum extraction and spectral modeling. XSNAP is designed to streamline all the processes utilized in Sections 2 and 3, e.g., source detection, deriving count-rates, spectral fitting and deriving unshocked CSM densities. The spectrum extraction component standardizes the procedures for each observatory described in Section 2, implementing them as dedicated Python Command-line Interface scripts. In addition, a script for generating source and background regions is provided, making use of ximage and (optionally) SAOImage DS9. For NuSTAR spectrum extraction process, we follow the standard guide from NuSTAR Data Analysis Software (NuSTARDAS [^4] https://heasarc.gsfc.nasa.gov/docs/nustar/analysis/nustar_quickstart_guide.pdf) (F. A. Harrison et al., 2013) inside HEASoft. Similarly, we developed the spectrum and CSM analysis sector by wrapping all the processes from Section 3 to each separate Python module. We then packaged the pipeline in a Python Package Index (PyPI) environment and published it to Github [^5] https://github.com/fercananything/XSNAP/ and PyPI [^6] https://xsnap.org/. Accordingly, users can streamline the whole analysis process by (i) making the source and background regions and (ii) extracting and reducing raw spectrum through the Python CLI scripts and (iii) analyzing spectrum from the Python API modules.
5 Discussion
Detected X-ray observations of SN 2024ggi up to $\sim 85$ days after explosion constrain both the shock properties and the progenitor activity before core collapse. SN 2024ggi has a progenitor mass-loss rate of $(6.2\pm 0.2)\times 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ for a steady-state wind velocity of $v_{\rm w}=20\ \rm km\ s^{-1}$ for the last 117 years before first light based on the phases of detected X-ray emission. This mass-loss rate is higher than what is estimated for most SNe IIP with X-ray observations (e.g., see I. Irani et al. 2024; V. V. Dwarkadas 2025), but comparable to the SN 2023ixf progenitor mass-loss rate of $\sim 10^{-4}\ \rm\ M_{\odot}\ yr^{-1}$ (A. J. Nayana et al., 2025; W. V. Jacobson-Galán et al., 2025a; K. A. Bostroem et al., 2025) at shock radii $10^{15}$ cm. Based on Figure 6 in J. Fuller & D. Tsuna (2024), $6.2\times 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ corresponds to a RSG progenitor mass of $\sim 20\ \rm M_{\odot}$. This estimate is higher than masses typically inferred from pre-explosion imaging and nebular-phase spectroscopy for CSM-interacting SNe II (e.g., C. D. Kilpatrick et al. 2025; W. V. Jacobson-Galán et al. 2025b), including SN 2024ggi (L. Ferrari et al., 2025; D. Xiang et al., 2024). Additionally, the X-ray luminosity peaks around the first $\sim 5$ days after first light and decays as $L_{X}\propto t^{-0.99}$, which is approximately aligned with a steady-state wind mass-loss who has a luminosity decay of $L\propto t^{-1}$.
A temporal evolution of the intrinsic neutral hydrogen column $(\rm NH_{\rm int})$ provides evidence for a dense, confined CSM (see Table 1). Although our results agree with W. V. Jacobson-Galán et al. (2024a) that a dense, confined CSM is present, the densities inferred from the optical spectra are higher than those derived from the X-rays at $5\times 10^{14}$ cm. This discrepancy, also observed in SN 2023ixf, may arise from CSM asymmetry or clumping (A. Singh et al., 2024; A. J. Nayana et al., 2025; S. S. Vasylyev et al., 2025). At early times $(\delta t\lesssim 16\ \rm days)$, we found $\rm NH_{\rm int}>10^{22}\ \rm cm^{-2}$, indicating significant X-ray absorption in a dense medium extending up to $r\lesssim 10^{15}\ \rm cm$, before this material overtaken by the SN shock at $\delta t\gtrsim 11\ \rm days$. All of these results were produced with our pipeline, XSNAP, which streamlines the entire workflow from raw counts to CSM analysis. Beyond this application to SN 2024ggi, XSNAP is broadly applicable: its data reduction and spectral modeling modules can be used for other X-ray sources, while its CSM analysis is specialized for SNe.
6 Conclusion
Using our XSNAP pipeline, we present our X-ray observation and analysis of SN 2024ggi, constraining both the shock properties and the mass loss rate of its RSG progenitor. In the following, we summarize our main findings of SN 2024ggi:
  • We estimate a progenitor mass-loss rate of $(6.2\pm 0.2)\times 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ from the detected emission for $\sim 117$ years before explosion $(v_{\rm wind}=20\ \rm km\ s^{-1})$.
  • SN 2024ggi shows evidence for a dense CSM $(\rm NH_{\rm int}>10^{22}\ \rm cm^{-2})$ up to $\sim 16$ days after explosion, confined within $r\lesssim 1.4\times 10^{15}\ \rm cm$.
  • SN 2024ggi has a luminosity decline of $L_{X}\propto t^{-0.99}$ and density profile $\rho\propto r^{-2}$, consistent with SN ejecta interaction in a steady-state, wind-like CSM.
  • All results were obtained with our new pipeline, XSNAP, demonstrating reproducibility from raw data reduction to CSM analysis. While specialized for supernovae, its reduction and spectral modeling modules are general and applicable to other X-ray sources. XSNAP is publicly available on GitHub and PyPI for community use.
SN 2024ggi offers a unique window into the study of SNe II and the characterization of presence of dense, confined CSM around their progenitor stars. Moreover, XSNAP enables a consistent framework to extend such studies to future X-ray–detected SNe.
$F.$ acknowledges support from the Caltech Summer Undergraduate Research Fellowship (SURF) program. W.J.-G. is supported by NASA through Hubble Fellowship grant HSTHF2-51558.001-A awarded by the Space Telescope Science Institute, which is operated for NASA by the Association of Universities for Research in Astronomy, Inc., under contract NAS5-26555. This research has made use of data and/or software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC.
Appendix A Contour Plot of Best Fit Parameter
Figure 4: Top left: Contour plot of the best fit parameter values for the column density and the Bremsstrahlung normalization from XRT data at $\delta t=3.33\ \rm days$. Top right: Contour plot of the best fit parameter values for the column density and the Bremsstrahlung normalization from XRT data at $\delta t=8.36\ \rm days$. Bottom left: Contour plot of the best fit parameter values for the column density and the Bremsstrahlung normalization from CXO data at $\delta t=10.90\ \rm days$. Bottom right: Contour plot of the best fit parameter values for the column density and the Bremsstrahlung normalization from CXO data at $\delta t=16.20\ \rm days$. Both parameters are well constrained in all the epochs, i.e. at $\delta t=3.33\ \rm days$, $\delta t=8.36\ \rm days$, $\delta t=10.90\ \rm days$, $\delta t=16.20\ \rm days$. The blue line represents $1\sigma$ range, the green line represents $2\sigma$ range, and the red line represents $3\sigma$ range.
Appendix B X-Ray Observation Logs
Table 2: X-ray Observation Log of SN2024ggi
Instrument Observation Date (yyyy/mm/dd) Mid Timea (days) Observation ID Exposure Time (ks) PI
Swift-XRT 2024-04-11 - 2024-04-16 3.33 00045607063, 00016601001, 00016601002, 00016601003, 00016601004, 00016601005, 00016601006, 00016601007, 00016601012, 00016601013, 00016601014, 00016601015, 00016601017, 00016601018, 00016601019, 00045607064, 00045607065 20 Swift, Sand
Swift-XRT 2024-04-17 - 2024-04-20 8.36 00016601020, 00016601022, 00016601023, 00016601024, 00016601025, 00016601026, 00016601027 9 Swift, Sand
CXO 2024-04-21 10.9 29383 14.9 E. Zimmerman
CXO 2024-04-26 16.2 29384 14.6 E. Zimmerman
Swift‐XRT 2024-05-01 – 2024-05-09 24.49 00016601044, 00016601045, 00016601048, 00016601049, 00016601050, 00016601051, 00016601052, 00016601053 7.6 D. Sand
Swift‐XRT 2024-05-12 – 2024-05-18 34.76 00016601055, 00016601059, 00016601062, 00016601063 5.0 D. Sand
Swift‐XRT 2024-05-21 – 2024-05-26 43.2 00016601064, 00016601065, 00016601066, 00016601067, 00016601068 10.1 D. Sand
XMM/EPIC-pn 2024-06-04 55.03 0882480901 16.9 W. Jacobson-Galan
XMM/EPIC-MOS1 2024-06-04 55.03 0882480901 16.2 W. Jacobson-Galan
XMM/EPIC-MOS2 2024-06-04 55.03 0882480901 16.2 W. Jacobson-Galan
Swift‐XRT 2024-06-02 – 2024-06-08 55.7 00016601069, 00016601070, 00016601071 2.4 D. Sand
Swift‐XRT 2024-06-24 – 2024-07-08 81.51 00016601072, 00016601073, 00016601074 6.1 D. Sand
XMM/EPIC-pn 2024-07-05 85.4 0882481001 9.8 W. Jacobson-Galan
XMM/EPIC-MOS1 2024-07-05 85.4 0882481001 11.6 W. Jacobson-Galan
XMM/EPIC-MOS2 2024-07-05 85.4 0882481001 11.6 W. Jacobson-Galan
Swift‐XRT 2024-07-15 – 2024-08-05 106.35 00016601075, 00016601076, 00016601077, 00016601078 4.7 D. Sand
Swift‐XRT 2024-08-08 – 2024-08-19 125.21 00016601079, 00016601080, 00016601081 4.9 D. Sand
Swift‐XRT 2024-11-04 208.13 00045607066 2.8 Swift
Swift‐XRT 2024-12-21 – 2025-01-01 260.33 00045607067, 00045607068, 00045607069, 00045607070 5.2 Swift
Swift‐XRT 2025-01-08 – 2025-01-15 276.65 00045607071, 00045607072, 00045607075 4.0 Swift
Swift‐XRT 2025-02-08 – 2025-03-05 316.43 00045607076, 00045607078, 00045607080 3.7 Swift
Swift‐XRT 2025-03-10 – 2025-03-20 338.76 00045607081, 00045607082 2.9 Swift
Notes: a With respect to time of explosion

### v2

@@
 XSNAP introduces application programming interfaces for per-epoch spectral modeling through PyXspec and emcee Markov chain Monte Carlo fitting.
-We employ XSNAP to model the multi-epoch X-ray spectra of SN 2024ggi with an absorbed thermal bremsstrahlung model and calculate a steady progenitor mass-loss rate of $(6.2\pm 0.2)\times 10^{-5}\,M_{\odot}\,\mathrm{yr^{-1}}$ $(v_{\rm wind}=20\ \rm km\ s^{-1})$, for which the detected X-ray emission traces the final 117 years before explosion.
+We employ XSNAP to model the multi-epoch X-ray spectra of SN 2024ggi with an absorbed thermal bremsstrahlung model and calculate a steady progenitor mass-loss rate of $(6.2\pm 0.2)\times 10^{-5}\,M_{\odot}\,\mathrm{yr^{-1}}$ $(v_{\rm wind}=20\ \rm km\ s^{-1})$, for which the detected X-ray emission traces the final $\sim 117$ years before explosion.
 The software is publicly available on GitHub, with a released package on the Python Package Index (PyPI).
@@
 show]ff10@illinois.edu
-1 Introduction
+I Introduction
 X-ray observations of supernovae (SNe) have become increasingly common in the last few decades.
@@
 Itagaki, 2023).
-It was first observed in at X-ray wavelengths with the Swift X-ray Telescope (XRT) at $\sim 1$ day after first light, which was then complimented with follow-up observations from the Chandra X-ray Observatory (CXO), XMM-Newton, and Nuclear Spectroscopic Telescope Array (NuSTAR) (W.
+It was first observed in at X-ray wavelengths with the Neil Gehrels Swift Observatory X-ray Telescope (Swift-XRT) at $\sim 1$ day after first light, which was then complimented with follow-up observations from the Chandra X-ray Observatory (CXO), XMM-Newton (XMM), and Nuclear Spectroscopic Telescope Array (NuSTAR) (W.
 Jacobson-Galán, 2025).
@@
 Hueichapán et al., 2025).
-In this work, we present and analyze the X-ray observations of SN 2024ggi from Chandra X-ray Observatory, Swift X-ray Telescope, and XMM-Newton.
+In this work, we present and analyze the X-ray observations of SN 2024ggi from CXO, Swift-XRT, and XMM.
 SN 2024ggi is located at $(\alpha=\rm 11^{h}18^{m}22.09^{s},\delta=-32^{\circ}50^{\prime}15.26^{\prime\prime})$ in the host galaxy NGC 3621.
@@
 Saha et al., 2006).
-We present multi-epoch X-ray observations of SN 2024ggi in Section 2.
-Our analysis of SN 2024ggi and its surrounding CSM are described in Section 3.
-In Section 4, we introduce the pipeline we have developed, XSNAP, and its design and availability.
-We discuss our findings in Section 5 and summarize them in Section 6.
-2 Observations
+In Section II, we describe the design, availability and applications of the “X-ray Supernova Analysis Pipeline (XSNAP).” In Section III, we present multi-epoch X-ray observations of SN 2024ggi, and our analysis of SN 2024ggi and its surrounding CSM are described in Section IV.
+We discuss our findings in Section V and summarize them in Section VI.
+II The Pipeline
+Each X-ray observatory, such as CXO, XMM, and Swift-XRT, and NuSTAR, employs mission-specific data formats, calibration pipelines, and analysis software.
+Consequently, this hinders reproducibility in analyzing X-ray observations as astronomers need to work out distinct guidelines for each mission.
+While there have been past efforts to streamline the process, such as YAXX [^1] https://cxc.harvard.edu/contrib/yaxx (Yet Another X-ray Xtractor; T.
+Aldcroft 2006), which supported both CXO and XMM-Newton data, these tools are no longer actively maintained.
+More recent softwares, like XGA [^2] https://xga.readthedocs.io/en/latest/ (X-ray: Generate and Analyze; D.
+J.
+Turner et al.
+2022) or BXA [^3] https://johannesbuchner.github.io/BXA/ (Bayesian X-ray Analysis; J.
+Buchner et al.
+2014), offer modern functionality but are limited to a single instrument (e.g., XMM), not designed as end-to-end pipelines from spectral extraction, or not specifically developed for SNe, even though some may still be adaptable for such studies.
+At present, there is no open-source, Python-based pipeline that unifies spectral extraction and analysis of SN X-ray spectra from CXO, XMM, Swift-XRT, and NuSTAR.
+Consequently, we developed XSNAP (X-ray Supernova Analysis Pipeline), an open-source Python package intended to unify end-to-end X-ray data processes from spectral extraction to CSM analysis.
+XSNAP is organized into two stages: spectrum extraction and spectral modeling and is designed to streamline all the processes described in Sections III and IV, including source detection, deriving count-rates, spectral fitting and deriving unshocked CSM densities.
+A typical XSNAP workflow consits of (i) generating source and background regions, (ii) extracting and reducing raw spectrum through the Python Command-line Interface (CLI) scripts, and (iii) analyzing spectrum from the Python Application Programming Interface (API) modules.
+This unified workflow is applied to the analysis of SN 2024ggi presented in Sections III and IV.
+The spectrum extraction component standardizes the procedures for observations from CXO, XMM, and Swift-XRT (Section III), implementing them as dedicated Python CLI scripts.
+In addition, a script is made seperately for observations from NuSTAR; we follow the standard guide from NuSTAR Data Analysis Software (NuSTARDAS [^4] https://heasarc.gsfc.nasa.gov/docs/nustar/analysis/nustar_quickstart_guide.pdf) (F.
+A.
+Harrison et al., 2013) inside HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014)Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), Heasarc).
+XSNAP also provides a script for generating source and background regions, making use of ximage and (optionally) SAOImage DS9 (W.
+A.
+Joye & E.
+Mandel, 2003).
+Similarly, we developed the spectrum and CSM analysis sector by wrapping all the processes from Section IV to each separate Python API module.
+We then packaged the pipeline in a Python Package Index (PyPI) environment and published it to Github [^5] https://github.com/fercananything/XSNAP/, PyPI [^6] https://pypi.org/p/xsnap/, and its own documentation website [^7] https://xsnap.org/.
+III Observations
 Table 1: Best‐fitting Parameters of the Absorbed Thermal Bremsstrahlung Model for SN2024ggi
 […]
-(days) $\left(10^{22}\ \rm cm^{-2}\right)$ (keV) (Absorbed) (Unabsorbed) ($10^{-4}$)
+(days) $\left(10^{22}\ \rm cm^{-2}\right)$ (keV) (Absorbed) (Unabsorbed) ($10^{-4}\ \rm\mathbf{cm^{-5}}$)
 3.33 Swift-XRT $7.36^{+3.93}_{-2.89}$ 47.81 $12.57^{+0.13}_{-0.14}$ $12.30^{+0.13}_{-0.14}$ $0.96^{+0.41}_{-0.30}$
@@
 Observed spectra are represented in blue (EPIC-pn), orange (EPIC-MOS1), and magenta (EPIC-MOS2) and black dashed lines are the best-fit absorbed thermal Bremsstrahlung model.
-Figure 2: Left: Plot of fitted unabsorbed $0.3-10\ \rm keV$ X-ray luminosity light curve from XMM-Newton (red circles), CXO (cyan squares), and Swift-XRT (orange triangles) observations.
+Figure 2: Left: Plot of fitted unabsorbed $0.3-10\ \rm keV$ X-ray luminosity light curve from XMM (red circles), CXO (cyan squares), and Swift-XRT (orange triangles) observations.
 The red dashed line is the fitted line with $L\propto t^{-0.99}$.
@@
 The grey dashed lines are reference density profile lines if the mass-loss rates are $10^{-4}\ \rm M_{\odot}\ yr^{-1}$, $10^{-5}\ \rm M_{\odot}\ yr^{-1}$, and $10^{-6}\ \rm M_{\odot}\ yr^{-1}$ respectively for $v_{\rm wind}=20\ \rm km\ s^{-1}$.
-2.1 Chandra X-ray Observatory
-The Chandra X-ray Observatory (CXO) ACIS-S observed SN 2024ggi in two epochs: one on 2024 April 21 (ObsID: 29383; $\delta t=10.90\ \rm days$; PI: E.
+III.1 Chandra X-ray Observatory
+The CXO ACIS-S observed SN 2024ggi in two epochs: one on 2024 April 21 (ObsID: 29383; $\delta t=10.90\ \rm days$; PI: E.
 Zimmerman) with an exposure time of 14.88 ks and the other on 2024 April 26 (ObsID: 29384; $\delta t=16.20\ \rm days$; PI: E.
@@
 Fruscione et al., 2006) and their corresponding calibration files.
-Source detection and estimating source-free background regions were utilized through ximage v4.5.1 from HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014)Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), Heasarc) with signal-to-noise ratio threshold of three $(\rm SNR_{thresh}=3)$ and human vetting with SAOImage DS9 (W.
-A.
-Joye & E.
-Mandel, 2003).
+Source detection and estimating source-free background regions were utilized through ximage v4.5.1 from HEASoft with signal-to-noise ratio threshold of three $(\rm SNR_{thresh}=3)$ and human vetting with SAOImage DS9.
 For each epoch, we extracted the spectrum via specextract and used a circular source region with a radius of $2^{\prime\prime}$ and circular background region with a radius of $45^{\prime\prime}$.
-2.2 XMM-Newton
-Similar to CXO, XMM-Newton (XMM) observed SN 2024ggi in two epochs: one on 2024 June 04 (ObsID: 0882480901; $\delta t=55.03\ \rm days$; PI: W.
+III.2 XMM-Newton
+Similar to CXO, XMM observed SN 2024ggi in two epochs: one on 2024 June 04 (ObsID: 0882480901; $\delta t=55.03\ \rm days$; PI: W.
 V.
-Jacobson-Gálan) and the other on 2024 July 5 (ObsID: 0882481001; $\delta t=85.40\ \rm days$; PI: W.
+Jacobson-Gálan) and the other on 2024 July 05 (ObsID: 0882481001; $\delta t=85.40\ \rm days$; PI: W.
 V.
@@
 European Photon Imaging Cameras (EPIC)-pn, MOS1, and MOS2, using the Scientific Analysis System (SAS) v22.1.0 (SAS development team, 2014) and corresponding calibration files.
-Source detection and estimating source-free background regions were utilized through ximage v4.5.1 from HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014)Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), Heasarc) with signal-to-noise ratio threshold of three $(\rm SNR_{thresh}=3)$ and human vetting with SAOImage DS9 (W.
-A.
-Joye & E.
-Mandel, 2003).
+Source detection and estimating source-free background regions were utilized through ximage v4.5.1 from HEASoft with signal-to-noise ratio threshold of three $(\rm SNR_{thresh}=3)$ and human vetting with SAOImage DS9.
 For each epoch, we extracted the spectrum via evselect and used a circular source region with a radius of $25^{\prime\prime}$ and circular background region with a radius of $125^{\prime\prime}$.
-2.3 Swift X-ray Telescope
-The X-Ray Telescope (XRT; D.
+III.3 Swift X-ray Telescope
+The Swift-XRT (D.
 N.
-Burrows et al.
-2005) onboard the Neil Gehrels Swift Observatory (N.
+Burrows et al., 2005) onboard the Neil Gehrels Swift Observatory (N.
 Gehrels et al., 2004) started observing SN 2024ggi on 2024 April 11 UT 14:04:10 $(\delta t=0.79\ \rm days)$.
-In this work, we analyzed XRT datasets collected until 2025 March 20.
+In this work, we analyzed Swift-XRT datasets collected until 2025 March 20.
 As each observation has relatively small exposure time $(\sim 1.5\ \rm ks)$, we binned a few epochs together, ranging from 10-day to 30-day intervals until the source was detected.
-However, SN 2024ggi was not detected in XRT images after 2024 June 02 $(\delta t=52.20\ \rm days)$ and these observations were analyzed as upper-limits (see Section 3).
+However, SN 2024ggi was not detected in Swift-XRT images after 2024 June 02 $(\delta t=52.20\ \rm days)$ and these observations were analyzed as upper-limits (see Section IV).
 Details of each stacked observation is reported in Table 2.
-We reduced each stacked observation using HEASoft v6.35 (Nasa High Energy Astrophysics Science Archive Research Center (2014)Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), Heasarc), the Swift XRT Data Analysis Software (SWXRTDAS; version 3.7.0), and the Swift Calibration Database (CALDB) version 20250609.
-Source detection and estimating source-free background regions were utilized through ximage v4.5.1 from HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014)Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), Heasarc) with signal-to-noise ratio threshold of three $(\rm SNR_{thresh}=3)$ and human vetting with SAOImage DS9 (W.
-A.
-Joye & E.
-Mandel, 2003).
+We reduced each stacked observation using HEASoft v6.35, the Swift XRT Data Analysis Software (SWXRTDAS; version 3.7.0), and the Swift Calibration Database (CALDB) version 20250609.
+Source detection and estimating source-free background regions were utilized through ximage v4.5.1 from HEASoft with signal-to-noise ratio threshold of three $(\rm SNR_{thresh}=3)$ and human vetting with SAOImage DS9.
 For each epoch, we extracted the spectrum via xrtpipeline, followed the standard practice from P.
@@
 (2012), and used a circular source region with a radius of $25^{\prime\prime}$ and circular background region with a radius of $125^{\prime\prime}$.
-3 Analysis
-3.1 X-ray Spectral Modeling
+IV Analysis
+IV.1 X-ray Spectral Modeling
 We modeled and fit each observation in Table 2 with PyXspec (C.
@@
 Arnaud, 1996).
-We fit them separately except for XMM-Newton data where we did a joint fit for all three instruments, EPIC-pn, MOS1, and MOS2.
-In this work, we adopt the solar abundances from M.
+We fit them separately except for XMM data where we did a joint fit for all three instruments, EPIC-pn, MOS1, and MOS2.
+In this work, we adopted the solar abundances from M.
 Asplund et al.
 (2009), i.e., abund aspl within PyXspec.
-We set the Galactic neutral hydrogen (NH) column density in the direction of SN2024ggi of $\mathrm{NH}_{\mathrm{MW}}=6.59\cdot 10^{20}\ \mathrm{cm}^{-2}$ (HI4PI Collaboration et al., 2016).
-For the spectral fits, we restricted the Chandra/ACIS data to the 0.58keV0.5-8\ $\mathrm${keV} energy range, while the Swift-XRT and XMM-Newton spectra were fit in the 0.310keV0.3-10\ $\mathrm${keV} band.
+We set the Galactic neutral hydrogen ($\rm N_{H}$) column density in the direction of SN2024ggi of $\mathrm{N}_{\mathrm{H,\ MW}}=6.59\cdot 10^{20}\ \mathrm{cm}^{-2}$ (HI4PI Collaboration et al., 2016).
+For the spectral fits, we restricted the Chandra/ACIS data to the 18keV1-8\ $\mathrm${keV} energy range, while the Swift-XRT and XMM spectra were fit in the 0.310keV0.3-10\ $\mathrm${keV} band.
 We then fit these data using an absorbed thermal Bremsstrahlung model (tbabs*ztbabs*bremss), but were unable to constrain the model temperature.
@@
 (2024a).
-Finally, we successfully modeled the intrinsic NH and Bremsstrahlung normalization and estimated the absorbed and unabsorbed fluxes in $0.3-10\ \rm keV$.
-We also repeat this analysis while adopting the temperature evolution of SN 2023ixf, $T\propto t^{-0.5}$ (A.
+Finally, we modeled the intrinsic $\rm N_{H}$ and Bremsstrahlung normalization and estimated the absorbed and unabsorbed fluxes in $0.3-10\ \rm keV$.
+We also repeated this analysis while adopting the temperature evolution of SN 2023ixf, $T\propto t^{-0.5}$ (A.
 J.
-Nayana et al., 2025), which yields consistent results to those found with a self-similar solution temperature decline rate.
+Nayana et al., 2025), which yielded consistent results to those found with a self-similar solution temperature decline rate.
 Based on the resulting best‐fit parameters, the X-ray spectra of SN 2024ggi reveal SN ejecta interaction with dense CSM: (i) the emission was highly absorbed at early times, (ii) the absorption decreased with time, and (iii) of thermal nature, consistent with previous studies (R.
@@
 The best-fit parameters are reported in Table 1 and contour plots of the first four epochs data from Swift-XRT and CXO are shown in 4.
-In addition, the fitted spectrum for CXO and XMM-Newton observations are shown in Figure 1.
-For observations where SN 2024ggi was not detected, we simulated the spectrum using the FakeIt command from PyXspec and used the faked spectrum flux and count rate ratio to estimate the flux upper-limit in $0.3-10\ \rm keV$.
-We set the intrinsic NH for non-detections to be $\rm NH_{int}=0$ and did not include their best-fit parameters for further analysis.
-We converted all the unabsorbed fluxes to luminosities and fit a power-law with emcee Markov chain Monte Carlo (MCMC).
+In addition, the fitted spectrum for CXO and XMM observations are shown in Figure 1.
+We classified an observation as a non-detection when the source, SN 2024ggi, has a signal-to-noise ratio below three $(\rm SNR<3)$.
+For those observations, we used the FakeIt command from PyXspec to simulate fake spectra with the same absorbed thermal Bremsstrahlung model (tbabs*ztbabs*bremss).
+In the model, we set the intrinsic $\rm N_{H}$ for non-detections to be $\rm N_{H,\ int}=0$ and fixed the temperature to follow $T=34\ \rm keV\ (t/13d)^{-0.25}$.
+We then used the fake spectrum flux and the count rate ratio to estimate the flux upper-limit in $0.3-10\ \rm keV$ band.
+However, we did not include parameters derived from non-detections for fitting the luminosity decay and the CSM density profile analysis.
+We converted all unabsorbed fluxes to luminosities and fit a power-law with emcee Markov chain Monte Carlo (MCMC).
 We found that the intrinsic $0.3-10\ \rm keV$ luminosity has a decay of $L_{X}\propto t^{-0.99}$ at $\delta t\gtrsim 8.36$ days.
@@
 (2025).
-3.2 CSM Density Profile Analysis
+IV.2 CSM Density Profile Analysis
 We estimated the CSM density profile, $\rho_{\rm CSM}(r)$, by using the normalization of thermal bremsstrahlung model as it is directly proportional to the emission measure (EM). EM is described by:
@@
 $\rm Bremss_{Norm}\equiv\frac{3.02\times 10^{-15}}{4\pi d^{2}}EM$ (2)
-where $n_{e}$ and $n_{I}$ are in $\rm cm^{-3}$ and $d$ is the distance to the source in $\rm cm^{-1}$.
+where $n_{e}$ and $n_{I}$ are in $\rm cm^{-3}$ and $d$ is the distance to the source in $\rm\mathbf{c}m$.
 We follow the procedure of D.
@@
 where $\mu_{e}$ and $\mu_{I}$ are the mean molecular weight of electron and ion respectively; $m_{p}$ is the proton mass; $\mathrm{V_{FS}}=\frac{4\pi}{3}f\left(R_{\rm out}^{3}-R_{\rm in}^{3}\right)$ is the emitting forward shock (FS) volume, with $R_{\rm in}$ and $R_{\rm out}$ are the inner and outer radius of the shell respectively and $f$ is the filling factor.
-In this work, we adopt $R_{\rm out}=1.2R_{\rm in}$, $f=1$, $\mu_{e}=1.14$, and $\mu_{I}=1.24$ (assuming solar values).
+In this work, we adopted $R_{\rm out}=1.2R_{\rm in}$, $f=1$, $\mu_{e}=1.14$, and $\mu_{I}=1.24$ (assuming solar values).
 We also estimated the shock radius by $R_{\rm shock}=v_{\rm shock}\times t$, where $t$ is time since explosion and assumed an average FS velocity $v_{\rm shock}=10^{4}\ \rm km\ s^{-1}$.
 The mass-loss rate was then estimated by fitting the density profile to $\rho_{CSM}(r)=\frac{\dot{M}}{4\pi r^{2}v_{\rm wind}}$ with the emcee package.
-Incorporating $v_{\rm wind}=20\ \rm km\ s^{-1}$, we inferred a mass-loss rate of $\dot{M}=(6.2\pm 0.2)\times 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ from the detected emission for the last 117 years before explosion, which lies between the upper-limit of the mass-loss rates of Type IIP SNe $\sim 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ and the lower-limit of the mass-loss rates of Type IIn SNe $\sim 10^{-3}\ \rm M_{\odot}\ yr^{-1}$ (V.
+Incorporating $v_{\rm wind}=20\ \rm km\ s^{-1}$, we inferred a mass-loss rate of $\dot{M}=(6.2\pm 0.2)\times 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ from the detected emission for the last $\sim 117$ years before explosion, which lies between the upper-limit of the mass-loss rates of Type IIP SNe $\sim 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ and the lower-limit of the mass-loss rates of Type IIn SNe $\sim 10^{-3}\ \rm M_{\odot}\ yr^{-1}$ (V.
 V.
 Dwarkadas, 2025).
+This lookback time can be treated as an approximation as they heavily rely on the wind and shock velocities, which both are assumed.
 The density profile and its comparison to SN 2023ixf and density profile models by W.
@@
 (2024a) are shown in Figure 3.
-4 The Pipeline
-Each X-ray observatory, such as CXO, XMM-Newton, and Swift-XRT, and Nuclear Spectroscopic Telescope Array (NuSTAR), employs mission-specific data formats, calibration pipelines, and analysis software.
-Consequently, this hinders reproducibility in analyzing X-ray observations as astronomers need to work out distinct guidelines for each mission.
-While there have been past efforts to streamline the process, such as YAXX [^1] https://cxc.harvard.edu/contrib/yaxx (Yet Another X-ray Xtractor; T.
-Aldcroft 2006), which supported both CXO and XMM-Newton data, these tools are no longer actively maintained.
-More recent softwares, like XGA [^2] https://xga.readthedocs.io/en/latest/ (X-ray: Generate and Analyse; D.
-J.
-Turner et al.
-2022) or BXA [^3] https://johannesbuchner.github.io/BXA/ (Bayesian X-ray Analysis; J.
-Buchner et al.
-2014), offer modern functionality but are limited to a single instrument (e.g., XMM-Newton), not designed as end-to-end pipelines from spectral extraction, or not specifically developed for SNe, even though some may still be adaptable for such studies.
-At present, there is no open-source, Python-based pipeline that unifies spectral extraction and analysis of SN X-ray spectra from CXO, XMM-Newton, Swift-XRT, and NuSTAR.
-Consequently, we developed XSNAP (X-ray Supernova Analysis Pipeline), an open-source Python package intended to unify end-to-end X-ray data processes from spectral extraction to CSM analysis.
-XSNAP is organized into two stages: spectrum extraction and spectral modeling.
-XSNAP is designed to streamline all the processes utilized in Sections 2 and 3, e.g., source detection, deriving count-rates, spectral fitting and deriving unshocked CSM densities.
-The spectrum extraction component standardizes the procedures for each observatory described in Section 2, implementing them as dedicated Python Command-line Interface scripts.
-In addition, a script for generating source and background regions is provided, making use of ximage and (optionally) SAOImage DS9.
-For NuSTAR spectrum extraction process, we follow the standard guide from NuSTAR Data Analysis Software (NuSTARDAS [^4] https://heasarc.gsfc.nasa.gov/docs/nustar/analysis/nustar_quickstart_guide.pdf) (F.
-A.
-Harrison et al., 2013) inside HEASoft.
-Similarly, we developed the spectrum and CSM analysis sector by wrapping all the processes from Section 3 to each separate Python module.
-We then packaged the pipeline in a Python Package Index (PyPI) environment and published it to Github [^5] https://github.com/fercananything/XSNAP/ and PyPI [^6] https://xsnap.org/.
-Accordingly, users can streamline the whole analysis process by (i) making the source and background regions and (ii) extracting and reducing raw spectrum through the Python CLI scripts and (iii) analyzing spectrum from the Python API modules.
-5 Discussion
+V Discussion
 Detected X-ray observations of SN 2024ggi up to $\sim 85$ days after explosion constrain both the shock properties and the progenitor activity before core collapse.
-SN 2024ggi has a progenitor mass-loss rate of $(6.2\pm 0.2)\times 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ for a steady-state wind velocity of $v_{\rm w}=20\ \rm km\ s^{-1}$ for the last 117 years before first light based on the phases of detected X-ray emission.
+SN 2024ggi has a progenitor mass-loss rate of $(6.2\pm 0.2)\times 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ for a steady-state wind velocity of $v_{\rm w}=20\ \rm km\ s^{-1}$ for the last $\sim 117$ years before first light based on the phases of detected X-ray emission.
 This mass-loss rate is higher than what is estimated for most SNe IIP with X-ray observations (e.g., see I.
@@
 Additionally, the X-ray luminosity peaks around the first $\sim 5$ days after first light and decays as $L_{X}\propto t^{-0.99}$, which is approximately aligned with a steady-state wind mass-loss who has a luminosity decay of $L\propto t^{-1}$.
-A temporal evolution of the intrinsic neutral hydrogen column $(\rm NH_{\rm int})$ provides evidence for a dense, confined CSM (see Table 1).
+A temporal evolution of the intrinsic neutral hydrogen column $(\rm N_{\rm H,\ int})$ provides evidence for a dense, confined CSM (see Table 1).
 Although our results agree with W.
@@
 Vasylyev et al., 2025).
-At early times $(\delta t\lesssim 16\ \rm days)$, we found $\rm NH_{\rm int}>10^{22}\ \rm cm^{-2}$, indicating significant X-ray absorption in a dense medium extending up to $r\lesssim 10^{15}\ \rm cm$, before this material overtaken by the SN shock at $\delta t\gtrsim 11\ \rm days$.
+At early times $(\delta t\lesssim 16\ \rm days)$, we found $\rm N_{\rm H,\ int}>10^{22}\ \rm cm^{-2}$, indicating significant X-ray absorption in a dense medium extending up to $r\lesssim 10^{15}\ \rm cm$, before this material overtaken by the SN shock at $\delta t\gtrsim 11\ \rm days$.
 All of these results were produced with our pipeline, XSNAP, which streamlines the entire workflow from raw counts to CSM analysis.
 Beyond this application to SN 2024ggi, XSNAP is broadly applicable: its data reduction and spectral modeling modules can be used for other X-ray sources, while its CSM analysis is specialized for SNe.
-6 Conclusion
+VI Conclusion
 Using our XSNAP pipeline, we present our X-ray observation and analysis of SN 2024ggi, constraining both the shock properties and the mass loss rate of its RSG progenitor.
@@
   • We estimate a progenitor mass-loss rate of $(6.2\pm 0.2)\times 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ from the detected emission for $\sim 117$ years before explosion $(v_{\rm wind}=20\ \rm km\ s^{-1})$.
-  • SN 2024ggi shows evidence for a dense CSM $(\rm NH_{\rm int}>10^{22}\ \rm cm^{-2})$ up to $\sim 16$ days after explosion, confined within $r\lesssim 1.4\times 10^{15}\ \rm cm$.
+  • SN 2024ggi shows evidence for a dense CSM $(\rm N_{\rm H,\ int}>10^{22}\ \rm cm^{-2})$ up to $\sim 16$ days after explosion, confined within $r\lesssim 1.4\times 10^{15}\ \rm cm$.
   • SN 2024ggi has a luminosity decline of $L_{X}\propto t^{-0.99}$ and density profile $\rho\propto r^{-2}$, consistent with SN ejecta interaction in a steady-state, wind-like CSM.
@@
 Moreover, XSNAP enables a consistent framework to extend such studies to future X-ray–detected SNe.
-$F.$ acknowledges support from the Caltech Summer Undergraduate Research Fellowship (SURF) program.
+F.
+acknowledges support from the Caltech Summer Undergraduate Research Fellowship (SURF) program.
 W.J.-G.
@@
 This research has made use of data and/or software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC.
+This paper employs a list of Chandra datasets, obtained by the Chandra X-ray Observatory, contained in the Chandra Data Collection  https://doi.org/10.25574/cdc.531 (catalog DOI: 10.25574/cdc.531).
 Appendix A Contour Plot of Best Fit Parameter
@@
 Appendix B X-Ray Observation Logs
+[…]
 Table 2: X-ray Observation Log of SN 2024ggi
-Instrument Observation Date (yyyy/mm/dd) Mid Timea (days) Observation ID Exposure Time (ks) PI
-Swift-XRT 2024-04-11 - 2024-04-16 3.33 00045607063, 00016601001, 00016601002, 00016601003, 00016601004, 00016601005, 00016601006, 00016601007, 00016601012, 00016601013, 00016601014, 00016601015, 00016601017, 00016601018, 00016601019, 00045607064, 00045607065 20 Swift, Sand
-Swift-XRT 2024-04-17 - 2024-04-20 8.36 00016601020, 00016601022, 00016601023, 00016601024, 00016601025, 00016601026, 00016601027 9 Swift, Sand
+Instrument Observation Mid Observation ID Exposure PI
+Date TimeaaWith respect to time of explosion. Time (ks)
+(yyyy/mm/$d$) (days)
+Swift-XRT 2024-04-11 – 3.33 00045607063, 00016601001, 00016601002, 20 Swift,
+2024-04-16 00016601003, 00016601004, 00016601005, D. Sand
+00016601006, 00016601007, 00016601012,
+00016601013, 00016601014, 00016601015,
+00016601017, 00016601018, 00016601019,
+00045607064, 00045607065
+Swift-XRT 2024-04-17 – 8.36 00016601020, 00016601022, 00016601023, 9 Swift,
+2024-04-20 00016601024, 00016601025, 00016601026, D. Sand
+00016601027
 CXO 2024-04-21 10.9 29383 14.9 E. Zimmerman
 CXO 2024-04-26 16.2 29384 14.6 E. Zimmerman
-Swift‐XRT 2024-05-01 – 2024-05-09 24.49 00016601044, 00016601045, 00016601048, 00016601049, 00016601050, 00016601051, 00016601052, 00016601053 7.6 D. Sand
-Swift‐XRT 2024-05-12 – 2024-05-18 34.76 00016601055, 00016601059, 00016601062, 00016601063 5.0 D. Sand
-Swift‐XRT 2024-05-21 – 2024-05-26 43.2 00016601064, 00016601065, 00016601066, 00016601067, 00016601068 10.1 D. Sand
+Swift‐XRT 2024-05-01 – 24.49 00016601044, 00016601045, 00016601048, 7.6 D. Sand
+2024-05-09 00016601049, 00016601050, 00016601051,
+00016601052, 00016601053
+Swift‐XRT 2024-05-12 – 34.76 00016601055, 00016601059, 00016601062, 5.0 D. Sand
+2024-05-18 00016601063
+Swift‐XRT 2024-05-21 – 43.2 00016601064, 00016601065, 00016601066, 10.1 D. Sand
+2024-05-26 00016601067, 00016601068
 XMM/EPIC-pn 2024-06-04 55.03 0882480901 16.9 W. Jacobson-Galan
@@
 XMM/EPIC-MOS2 2024-06-04 55.03 0882480901 16.2 W. Jacobson-Galan
-Swift‐XRT 2024-06-02 – 2024-06-08 55.7 00016601069, 00016601070, 00016601071 2.4 D. Sand
-Swift‐XRT 2024-06-24 – 2024-07-08 81.51 00016601072, 00016601073, 00016601074 6.1 D. Sand
+Swift‐XRT 2024-06-02 – 55.7 00016601069, 00016601070, 00016601071 2.4 D. Sand
+2024-06-08
+Swift‐XRT 2024-06-24 – 81.51 00016601072, 00016601073, 00016601074 6.1 D. Sand
+2024-07-08
 XMM/EPIC-pn 2024-07-05 85.4 0882481001 9.8 W. Jacobson-Galan
@@
 XMM/EPIC-MOS2 2024-07-05 85.4 0882481001 11.6 W. Jacobson-Galan
-Swift‐XRT 2024-07-15 – 2024-08-05 106.35 00016601075, 00016601076, 00016601077, 00016601078 4.7 D. Sand
-Swift‐XRT 2024-08-08 – 2024-08-19 125.21 00016601079, 00016601080, 00016601081 4.9 D. Sand
+Swift‐XRT 2024-07-15 – 106.35 00016601075, 00016601076, 00016601077, 4.7 D. Sand
+2024-08-05 00016601078
+Swift‐XRT 2024-08-08 – 125.21 00016601079, 00016601080, 00016601081 4.9 D. Sand
+2024-08-19
 Swift‐XRT 2024-11-04 208.13 00045607066 2.8 Swift
-Swift‐XRT 2024-12-21 – 2025-01-01 260.33 00045607067, 00045607068, 00045607069, 00045607070 5.2 Swift
-Swift‐XRT 2025-01-08 – 2025-01-15 276.65 00045607071, 00045607072, 00045607075 4.0 Swift
-Swift‐XRT 2025-02-08 – 2025-03-05 316.43 00045607076, 00045607078, 00045607080 3.7 Swift
-Swift‐XRT 2025-03-10 – 2025-03-20 338.76 00045607081, 00045607082 2.9 Swift
-Notes: a With respect to time of explosion
+Swift‐XRT 2024-12-21 – 260.33 00045607067, 00045607068, 00045607069, 5.2 Swift
+2025-01-01 00045607070
+Swift‐XRT 2025-01-08 – 276.65 00045607071, 00045607072, 00045607075 4.0 Swift
+2025-01-15
+Swift‐XRT 2025-02-08 – 316.43 00045607076, 00045607078, 00045607080 3.7 Swift
+2025-03-05
+Swift‐XRT 2025-03-10 – 338.76 00045607081, 00045607082 2.9 Swift
+2025-03-20